# Minimax Hypothesis Testing for the Bradley-Terry-Luce Model

### v1

@@ IV-D Proof of Theorem 1
Note that $R$ is symmetric and is in fact a positive semidefinite matrix (as $u^{\mathrm{T}}Ru=\sum_{(i,j)\in{\mathcal{E}}}(u_{i}p_{ij}-u_{j}p_{ji})^{2}$). Therefore, all its eigenvalues are non-negative. Moreover, when the canonical Markov matrix corresponding to $P$ is not reversible, $R$ is, in fact, positive-definite. Additionally, when the canonical Markov matrix corresponding to $P$ is reversible, then Eq. 33 trivially holds as both sides of the inequality are zero. Therefore, our main focus will be on the case when the matrix $P$ is not reversible. First, to begin, let us focus on the denominator term $\pi_{B}^{*^{\mathrm{T}}}R\pi_{B}^{*}$. Recall that ${\pi_{B}^{*}}$ is the solution of the following optimization problem:
$\min_{\begin{subarray}{c}\pi_{B}\in{\mathbb{R}}^{n}:\pi_{B}^{\mathrm{T}}{\mathbf{1}}_{n}=1\end{subarray}}\frac{1}{2}\pi_{B}^{\mathrm{T}}R\pi_{B}.$
Forming the Lagrangian $L(\pi_{B},\lambda)$, we get
$\min_{\pi_{B}}L(\pi_{B},\lambda)\triangleq\frac{1}{2}\pi_{B}^{\mathrm{T}}R{\pi_{B}}-\lambda\left(\pi_{B}^{\mathrm{T}}{\mathbf{1}}_{n}-1\right).$
This gives the optimality condition as $R\pi_{B}^{*}=\lambda^{*}{\mathbf{1}}_{n}$, where $\lambda^{*}$ is the optimal dual solution. The optimality condition implies
@@
Therefore, $\lambda^{*}\geq 0$ as $R$ is positive (semi-)definite. Equivalently, the optimality condition implies that $\pi_{B}^{*}$ is the Perron-Frobenius eigenvector of the matrix $S_{B,\lambda^{*}}$ where $S_{B,\lambda}$ is a matrix with parameter $\lambda$ and is defined as
$S_{B,\lambda}\triangleq I_{n}+\frac{1}{n}(\lambda{\mathbf{1}}_{n}{\mathbf{1}}_{n}^{\mathrm{T}}-R).$ (37)
This is because the optimality condition implies that $\pi_{B}^{*^{\mathrm{T}}}=\pi_{B}^{*^{\mathrm{T}}}(I_{n}+\frac{\lambda^{*}{\mathbf{1}}_{n}{\mathbf{1}}_{n}^{\mathrm{T}}-R}{n})$. Moreover, note that each entry of $S_{B,\lambda}\geq 0$ for $\lambda\geq 0$. Now, consider the numerator term in Eq. 34, which can be expressed as
@@
where $\lambda_{\max}(R)$ is maximum eigenvalue of $R$. Note that by Gershgorin circle theorem [78], we have $\lambda_{\max}(R)\leq\max_{i}\sum_{j=1}^{n}|R_{ij}|\leq 2d_{\max}$. Now, we will upper bound $\|\pi-{\pi_{B}^{*}}\|_{2}$. Since, $\pi$ and ${\pi_{B}^{*}}$ are Perron vectors of $S$ and $S_{B,\lambda^{*}}$ with eigenvalues 1, therefore we have
$\pi^{\mathrm{T}}-\pi_{B}^{*^{\mathrm{T}}}=$ $\pi^{\mathrm{T}}S-\pi_{B}^{*^{\mathrm{T}}}S_{B,\lambda^{*}}=\left(\pi-\pi_{B}^{*}\right)^{\mathrm{T}}\left(S-{\mathbf{1}}_{n}\pi^{\mathrm{T}}\right)+\pi_{B}^{*^{\mathrm{T}}}(S-S_{B,\lambda^{*}})$
@@
$\|\pi-\pi_{B}^{*}\|_{\pi^{-1}}\leq\|\pi-\pi_{B}^{*}\|_{\pi^{-1}}\|S-{\mathbf{1}}_{n}\pi^{\mathrm{T}}\|_{\pi^{-1}}+\|\pi_{B}^{*^{\mathrm{T}}}(S-S_{B,\lambda^{*}})\|_{\pi^{-1}}.$ (39)
It is straightforward to verify that
@@
$\|\pi-{\pi_{B}^{*}}\|_{2}$ $\leq 4\sqrt{\frac{\|\pi\|_{\infty}}{\pi_{\min}}}\frac{\|{\pi_{B}^{*}}^{\mathrm{T}}(S-S_{B,\lambda^{*}})\|_{2}}{\xi^{2}}\leq\frac{4\sqrt{h}}{\xi^{2}}(\|{\pi_{B}^{*}}^{\mathrm{T}}(S-S_{B,0})\|_{2}+\frac{\lambda^{*}}{n}\|{\pi_{B}^{*}}^{\mathrm{T}}{\mathbf{1}}_{n}{\mathbf{1}}_{n}^{\mathrm{T}}\|_{2})$
$\leq O\bigg{(}\|{\pi_{B}^{*}}^{\mathrm{T}}(S-S_{B,0})\|_{2}+\frac{\lambda^{*}}{\sqrt{n}}\bigg{)}$ (40)
@@
$({\pi_{B}^{*}}^{\mathrm{T}}(S-S_{B,0}))_{i}$ $=\frac{1}{d}\bigg{(}\pi_{B,i}^{*}\sum_{\begin{subarray}{c}j:j\neq i\\ (i,j)\in{\mathcal{E}}\end{subarray}}\left(\frac{d}{n}p_{ij}^{2}-p_{ij}\right)+\sum_{\begin{subarray}{c}j:j\neq i\\ (i,j)\in{\mathcal{E}}\end{subarray}}\pi_{B,j}^{*}\left(p_{ji}-\frac{d}{n}p_{ij}p_{ji}\right)\bigg{)}$
$=-\frac{1}{d}\sum_{\begin{subarray}{c}j:j\neq i\\ (i,j)\in{\mathcal{E}}\end{subarray}}(\pi_{B,i}^{*}p_{ij}-\pi_{B,j}^{*}p_{ji})\left(1-\frac{d}{n}p_{ij}\right).$
@@
$\|{\pi_{B}^{*}}^{\mathrm{T}}(S-S_{B,0})\|_{2}^{2}$ $\leq\frac{1}{d^{2}}\sum_{i=1}^{n}\bigg{(}\sum_{\begin{subarray}{c}j:j\neq i\\ (i,j)\in{\mathcal{E}}\end{subarray}}|\pi_{B,i}^{*}p_{ij}-\pi_{B,j}^{*}p_{ji}|\bigg{)}^{2}\leq\frac{d_{\max}}{d^{2}}\sum_{(i,j)\in{\mathcal{E}}}(\pi_{B,i}^{*}p_{ij}-\pi_{B,j}^{*}p_{ji})^{2}$
@@
Combining Eqs. 40 and 41, we get
$\|\pi-{\pi_{B}^{*}}\|_{2}\leq O\bigg{(}\frac{\lambda^{*}}{\sqrt{n}}+\sqrt{\frac{\lambda^{*}}{d}}\bigg{)}.$
@@
$\pi^{\mathrm{T}}R\pi\leq\lambda^{*}+2{c}\lambda^{*}d_{\max}\bigg{(}\frac{1}{\sqrt{n}}+\sqrt{\frac{\lambda^{*}}{d}}\bigg{)}^{2}\leq\lambda^{*}+2c\lambda^{*}(1+\sqrt{\lambda^{*}/2})^{2}.$ (42)
Therefore, using Eqs. 36 and 42, the ratio in Eq. 34 is upper bounded as
$\frac{\|\Pi P-P^{\mathrm{T}}\Pi\|_{\mathrm{F}}}{\|{\Pi_{B}^{*}}P-P^{\mathrm{T}}{\Pi_{B}^{*}}\|_{\mathrm{F}}}\leq 1+O(\sqrt{\lambda^{*}}).$ (43)
Finally, it remains to show that $\lambda^{*}$ is upper bounded by a constant. We show this through the following argument. Note that $S_{B,0}$ is a symmetric matrix with absolute eigenvalues strictly less than $1$ (when the canonical Markov matrix corresponding to $P$ is not reversible). This is because for any vector $u\in{\mathbb{R}}^{n}$ with $\|u\|_{2}=1$, we have
@@
$\lambda_{\max}(S_{B,\lambda_{0}})\geq\min_{i\in[n]}\bigg{(}1+\lambda_{0}+\frac{1}{n}\sum_{\begin{subarray}{c}j:j\neq i\\ (i,j)\in{\mathcal{E}}\end{subarray}}(p_{ij}p_{ji}-p_{ij}^{2})\bigg{)}\geq\lambda_{0}+1-\frac{d_{\max}}{n}.$
Setting $\lambda_{0}=\frac{d_{\max}}{n}$ ensures $\lambda_{\max}(S_{B,\lambda_{0}})\geq 1$. Thus, $\lambda^{*}\leq\frac{d_{\max}}{n}\leq 1$, proving the theorem.
@@ V-A Preliminaries
In this section, we will prove key lemmata that will be used quite frequently to develop the proof of the main result in Theorem 2. The following lemma is similar to [11, Theorem 8] but also holds when the canonical Markov matrix of $P$ is not reversible.
@@
Let $\pi,{\hat{\pi}},\tilde{\pi}$ be the stationary distributions of the row stochastic matrices $S,{\hat{S}},\tilde{S}$, respectively such that $\pi>0$ entrywise. Then, if $\|\Pi^{1/2}S\Pi^{-1/2}-\sqrt{\pi}\sqrt{\pi}^{\mathrm{T}}\|_{2}+\|S-\hat{S}\|_{\pi^{-1}}<1$, we have
$\|{\hat{\pi}}-\tilde{\pi}\|_{\pi^{-1}}\leq\frac{\|\tilde{\pi}^{\mathrm{T}}(\tilde{S}-{\hat{S}})\|_{\pi^{-1}}}{1-\|\Pi^{1/2}S\Pi^{-1/2}-\sqrt{\pi}\sqrt{\pi}^{\mathrm{T}}\|_{2}-\|S-\hat{S}\|_{\pi^{-1}}}.$
Proof: By stationarity of $\tilde{S}$ and ${\hat{S}}$, we have
@@
$\|\tilde{\pi}-{\hat{\pi}}\|_{\pi^{-1}}$ $\leq\|\tilde{\pi}^{\mathrm{T}}(\tilde{S}-{\hat{S}})\|_{\pi^{-1}}+\|\tilde{\pi}-{\hat{\pi}}\|_{\pi^{-1}}\|\hat{S}-S\|_{\pi^{-1}}+\|\tilde{\pi}-{\hat{\pi}}\|_{\pi^{-1}}\|S-\mathbf{1}_{n}\pi^{\mathrm{T}}\|_{\pi^{-1}}.$ (45)
@@
$\|S-\mathbf{1}_{n}\pi^{\mathrm{T}}\|_{\pi^{-1}}=\|\Pi^{1/2}S\Pi^{-1/2}-\sqrt{\pi}\sqrt{\pi}^{\mathrm{T}}\|_{2}=\sigma_{2}(\Pi^{1/2}S\Pi^{-1/2}),$
where $\sigma_{2}(M)$ denotes the second largest singular value of $M$ and the last equality follows since $\sqrt{\pi}$ is both the left and right top singular vector of the DTM $\Pi^{1/2}S\Pi^{-1/2}$. Thus, rearranging the terms in the above inequality establishes the statement of Lemma 1.
We would like to emphasize the distinction between our proof above and the approach adopted in [11, Theorem 8] for proving a similar result, and elucidate the proof in [11]. In transitioning from Eq. 44 to Eq. 45, the authors in [11] utilize the $\ell^{2}(\pi)$ norm, as opposed to the $\ell^{2}(\pi^{-1})$ norm. Specifically, they bound the final term on the right-hand side as follows
@@
Moreover, since the Markov chain is irreducible by the Perron-Frobenius theorem, we have $\lambda_{2}(S)<1$, and the corresponding upper bounds have been derived in [30, 11]. However, we are interested in the general case where $S$ need not be reversible (as the underlying pairwise comparison matrix may not be BTL). Hence, we bound each of these terms using the following two lemmata.
Lemma 2 (Spectral Norm of Noise [30]).
For ${\hat{S}}$ constructed as in Eq. 25, we have
@@
The proof of Lemma 2 is quite similar to [30] but requires slight modifications and is provided in Section V-C3 for completeness.
Next we upper bound the quantity $\|R-\sqrt{\pi}\sqrt{\pi}^{\mathrm{T}}\|_{2}$. Observe that $\|R-\sqrt{\pi}\sqrt{\pi}^{\mathrm{T}}\|_{2}$ is the second largest singular value of the DTM $R$ (as $\sqrt{\pi}$ are both the left and right singular-vectors of $R$ corresponding to largest singular-value of $1$ [58, Proposition 2.2]). The second largest singular value of DTM $R$ is also equal to the square root of the contraction coefficient for $\chi^{2}$-divergence $\eta_{\chi^{2}}(\pi,{S})$ of the source-channel pair $(\pi,S)$ (see [79, 58, 59] for definitions and details). For the case of the complete graph, we can upper bound this quantity by upper bounding the Dobrushin contraction coefficient for TV distance, as demonstrated in the following lemma.
@@
Proof: We will find an upper bound on $\sqrt{\eta_{\chi^{2}}(\pi,{S})}$ using [58, Proposition 2.5]:
$\eta_{\chi^{2}}(\pi,S)\leq\eta_{\mathsf{TV}}(S)=\max_{i,j}\ \left\|S_{i,:}-S_{j,:}\right\|_{\mathsf{TV}},$
@@
$\left\|{S}_{i,:}-{S}_{j,:}\right\|_{\mathsf{TV}}$ $=1-\sum_{k=1}^{n}\min\{S_{ik},S_{jk}\}=1-\left(\sum_{k:k\neq i,j}\frac{\min\{p_{ik},p_{jk}\}}{2n}+\frac{p_{ij}+p_{ji}}{2n}\right)$
@@
Hence, the lemma holds since for $0\leq x\leq 1$, we have $\sqrt{1-x}\leq 1-\frac{x}{2}$.
In the more general case of an arbitrary graph (consistent with our assumptions), we upper-bound the second largest singular value of $R$ by leveraging the edge expansion properties and Cheeger’s inequality for non-negative matrices as demonstrated in the following lemma:
@@
Let ${\mathcal{G}}$ be the induced graph corresponding to the canonical Markov matrix $S$. Consider the DTM $R=\Pi^{1/2}S\Pi^{-1/2}$, with edge expansion lower bounded by $\xi$, i.e., $\phi(R)\geq\xi$, then the following bounds hold on the second largest singular value of $R$:
@@
Proof: Observe that
$\sigma_{2}(R)$ $=\sqrt{\lambda_{2}\left(RR^{\mathrm{T}}\right)}\stackrel{{\scriptstyle\zeta_{1}}}{{\leq}}\sqrt{\lambda_{2}\left(\frac{R+R^{\mathrm{T}}}{2}\right)}\stackrel{{\scriptstyle\zeta_{2}}}{{\leq}}\sqrt{1-\frac{1}{2}\phi^{2}\left(\frac{R+R^{\mathrm{T}}}{2}\right)}$

### v2

@@ IV-D Proof of Theorem 1
Note that $R$ is symmetric and is in fact a positive semidefinite matrix (as $u^{\mathrm{T}}Ru=\sum_{(i,j)\in{\mathcal{E}}}(u_{i}p_{ij}-u_{j}p_{ji})^{2}$). Therefore, all its eigenvalues are non-negative. Moreover, when the canonical Markov matrix corresponding to $P$ is not reversible, $R$ is, in fact, positive definite. The positive definiteness of $R$ follows from the fact that if $u^{\mathrm{T}}Ru=0$ for some vector $u$, then $u_{i}p_{ij}=u_{j}p_{ji}$ for all $(i,j)\in\mathcal{E}$. By the connectedness of the graph $\mathcal{G}$}, this implies that $u\geq 0$ entrywise (without loss of generality). Consequently, the canonical Markov matrix $S$ corresponding to $P$ is reversible, which leads to a contradiction. Additionally, when the canonical Markov matrix corresponding to $P$ is reversible, then Eq. 31 trivially holds as both sides of the inequality are zero. Therefore, our main focus will be on the case when the matrix $P$ is not reversible. First, to begin, let us focus on the denominator term $\pi_{B}^{*^{\mathrm{T}}}R\pi_{B}^{*}$. Recall that $\pi_{B}^{*}$} is the solution of the following optimization problem:
$\min_{\begin{subarray}{c}\pi_{B}\in{\mathbb{R}}_{+}^{n}:\pi_{B}^{\mathrm{T}}{\mathbf{1}}_{n}=1\end{subarray}}\frac{1}{2}\pi_{B}^{\mathrm{T}}R\pi_{B}.$
For $u\geq{\mathbf{0}}$ and $\lambda\in{\mathbb{R}},$ we obtain the Lagrangian
$L(\pi_{B},\lambda,u)\triangleq\frac{1}{2}\pi_{B}^{\mathrm{T}}R{\pi_{B}}-\lambda\left(\pi_{B}^{\mathrm{T}}{\mathbf{1}}_{n}-1\right)-u^{\mathrm{T}}\pi_{B}.$
Using Karush-Kuhn-Tucker conditions, this gives the following optimality conditions (primal feasibility, dual feasibility, stationarity, and complementary slackness):
$\pi_{B}^{*}\geq 0,\ \pi_{B}^{*^{\mathrm{T}}}{\mathbf{1}}_{n}=1,\ u^{*}\geq{\mathbf{0}},\ R\pi_{B}^{*}-\lambda^{*}{\mathbf{1}}_{n}=u^{*},$
$\text{and}\ \forall i\in[n],\,u_{i}^{*}(\pi_{B}^{*})_{i}=0,$
where $\lambda^{*},u^{*}$ is the optimal dual solution. We claim that when $R$ is positive definite, the primal and dual solution satisfying the above optimality conditions are given by:
$\pi_{B}^{*}=\frac{R^{-1}{\mathbf{1}}_{n}}{{\mathbf{1}}_{n}^{\mathrm{T}}R^{-1}{\mathbf{1}}_{n}},\lambda^{*}=\frac{1}{{\mathbf{1}}_{n}^{\mathrm{T}}R^{-1}{\mathbf{1}}_{n}},u^{*}={\mathbf{0}}.$
Here, $\pi_{B}^{*}$ is non-negative, because $R^{-1}$ is entrywise non-negative. To see this latter fact, we use the inverse-positivity property of $R$, because $R$ is a special kind of $Z$-matrix (i.e., $R_{ij}\leq 0$ for $i\neq j$), namely, an $M$-matrix [ZmatrixMmatrix]. Moreover, the optimality conditions imply that
@@
Equivalently, the optimality condition implies that $\pi_{B}^{*}$ is an eigenvector of the matrix $S_{B,\lambda^{*}}$ with eigenvalue $1$, where $S_{B,\lambda}$ is a matrix with parameter $\lambda$ and is defined as
$S_{B,\lambda}\triangleq I_{n}+\frac{1}{n}(\lambda{\mathbf{1}}_{n}{\mathbf{1}}_{n}^{\mathrm{T}}-R),$ (35)
where $I_{n}$ is the identity matrix of size $n$. This is because the optimality condition implies that ${\pi_{B}^{*}}^{\mathrm{T}}={\pi_{B}^{*}}^{\mathrm{T}}\big(I_{n}+\frac{\lambda^{*}{\mathbf{1}}_{n}{\mathbf{1}}_{n}^{\mathrm{T}}-R}{n}\big)$. Moreover, note that each entry of $S_{B,\lambda^{*}}$ is strictly greater than 0 for $\lambda^{*}>0$, and therefore, $\pi_{B}^{*}$ is the Perron vector of $S_{B,\lambda^{*}}$ with eigenvalue 1. Now, consider the numerator term in Eq. 32, which can be expressed as
@@
where $\lambda_{\max}(R)$ is maximum eigenvalue of $R$. Note that by Gershgorin circle theorem [horn2012matrix], we have $\lambda_{\max}(R)\leq\max_{i}\sum_{j=1}^{n}|R_{ij}|\leq 2d_{\max}$. Now, we will upper bound $\|\pi-{\pi_{B}^{*}}\|_{2}$. Since, $\pi$ and $\pi_{B}^{*}$} are Perron vectors of $S$ and $S_{B,\lambda^{*}}$ with eigenvalues 1, therefore we have
$\pi^{\mathrm{T}}-\pi_{B}^{*^{\mathrm{T}}}=$ $\pi^{\mathrm{T}}S-\pi_{B}^{*^{\mathrm{T}}}S_{B,\lambda^{*}}$
$=\left(\pi-\pi_{B}^{*}\right)^{\mathrm{T}}\left(S-{\mathbf{1}}_{n}\pi^{\mathrm{T}}\right)+\pi_{B}^{*^{\mathrm{T}}}(S-S_{B,\lambda^{*}})$
@@
$\|\pi-\pi_{B}^{*}\|_{\pi^{-1}}$ $\leq\|\pi-\pi_{B}^{*}\|_{\pi^{-1}}\|S-{\mathbf{1}}_{n}\pi^{\mathrm{T}}\|_{\pi^{-1}}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\|\pi_{B}^{*^{\mathrm{T}}}(S-S_{B,\lambda^{*}})\|_{\pi^{-1}}.$ (37)
It is straightforward to verify that (see Section V-A for more details)
@@
$\|\pi\!-\!{\pi_{B}^{*}}\|_{2}$ $\leq 4\sqrt{\frac{\|\pi\|_{\infty}}{\pi_{\min}}}\frac{\|{\pi_{B}^{*}}^{\mathrm{T}}(S-S_{B,\lambda^{*}})\|_{2}}{\xi^{2}}$
$\leq\frac{4\sqrt{h}}{\xi^{2}}\bigg(\!\|{\pi_{B}^{*}}^{\mathrm{T}}(S-S_{B,0})\|_{2}+\frac{\lambda^{*}}{n}\|{\pi_{B}^{*}}^{\mathrm{T}}{\mathbf{1}}_{n}{\mathbf{1}}_{n}^{\mathrm{T}}\|_{2}\!\bigg)$
$\leq\frac{4\sqrt{h}}{\xi^{2}}\bigg(\|{\pi_{B}^{*}}^{\mathrm{T}}(S-S_{B,0})\|_{2}+\frac{\lambda^{*}}{\sqrt{n}}\bigg),$ (38)
@@
$({\pi_{B}^{*}}^{\mathrm{T}}(S-S_{B,0}))_{i}=\frac{1}{d}\bigg(\pi_{B,i}^{*}\sum_{\begin{subarray}{c}j:j\neq i,\\ (i,j)\in{\mathcal{E}}\end{subarray}}\left(\frac{d}{n}p_{ij}^{2}-p_{ij}\right)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\sum_{\begin{subarray}{c}j:j\neq i,\\ (i,j)\in{\mathcal{E}}\end{subarray}}\pi_{B,j}^{*}\left(p_{ji}-\frac{d}{n}p_{ij}p_{ji}\right)\bigg)$
$=-\frac{1}{d}\sum_{\begin{subarray}{c}j:j\neq i,\\ (i,j)\in{\mathcal{E}}\end{subarray}}(\pi_{B,i}^{*}p_{ij}-\pi_{B,j}^{*}p_{ji})\left(1-\frac{d}{n}p_{ij}\right).$
@@
$\|{\pi_{B}^{*}}^{\mathrm{T}}(S-S_{B,0})\|_{2}^{2}$ $\leq\frac{1}{d^{2}}\sum_{i=1}^{n}\bigg(\sum_{\begin{subarray}{c}j:j\neq i,\\ (i,j)\in{\mathcal{E}}\end{subarray}}|\pi_{B,i}^{*}p_{ij}-\pi_{B,j}^{*}p_{ji}|\bigg)^{2}$
$\leq\frac{d_{\max}}{d^{2}}\sum_{(i,j)\in{\mathcal{E}}}(\pi_{B,i}^{*}p_{ij}-\pi_{B,j}^{*}p_{ji})^{2}$
@@
Combining Eqs. 38 and 39, we get
$\|\pi-{\pi_{B}^{*}}\|_{2}\leq\frac{4\sqrt{h}}{\xi^{2}}\bigg(\frac{\lambda^{*}}{\sqrt{n}}+\sqrt{\frac{\lambda^{*}}{2d}}\bigg).$
@@
$\pi^{\mathrm{T}}R\pi$ $\leq\lambda^{*}+\frac{ch}{\xi^{4}}\lambda^{*}d_{\max}\bigg(\frac{1}{\sqrt{d}}+\sqrt{\frac{\lambda^{*}}{n}}\bigg)^{2}$
$\leq\lambda^{*}+\frac{ch}{\xi^{4}}\lambda^{*}\left(1+\sqrt{\lambda^{*}}\right)^{2}.$ (40)
Therefore, using Eqs. 34 and 40, the ratio in Eq. 32 is upper bounded as
$\frac{\|\Pi P-P^{\mathrm{T}}\Pi\|_{\mathrm{F}}}{\|{\Pi_{B}^{*}}P-P^{\mathrm{T}}{\Pi_{B}^{*}}\|_{\mathrm{F}}}\leq 1+O\bigg(\frac{\sqrt{h}}{\xi^{2}}\sqrt{\lambda^{*}}\bigg).$ (41)
Finally, it remains to show that $\lambda^{*}=1/{\mathbf{1}}_{n}^{\mathrm{T}}R^{-1}{\mathbf{1}}_{n}$ is upper bounded by a constant. We can show this using a spectral bound $\lambda^{*}\leq\lambda_{\max}(R)/n\leq 2d_{\max}/n\leq 2$.
Alternatively, we can also show this through the following argument. Note that $S_{B,0}$ is a symmetric matrix with absolute eigenvalues strictly less than $1$ (when the canonical Markov matrix corresponding to $P$ is not reversible). This is because for any vector $u\in{\mathbb{R}}^{n}$ with $\|u\|_{2}=1$, we have
@@
$\lambda_{\max}(S_{B,\lambda_{0}})$ $\geq\min_{i\in[n]}\bigg(1+\lambda_{0}+\frac{1}{n}\sum_{\begin{subarray}{c}j:j\neq i,\\ (i,j)\in{\mathcal{E}}\end{subarray}}(p_{ij}p_{ji}-p_{ij}^{2})\bigg)$
$\geq\lambda_{0}+1-\frac{d_{\max}}{n}.$
Setting $\lambda_{0}=\frac{d_{\max}}{n}$ ensures $\lambda_{\max}(S_{B,\lambda_{0}})\geq 1$. Thus, $\lambda^{*}\leq\frac{d_{\max}}{n}\leq 1$, proving the theorem. ∎
@@ V-A Preliminaries
In this section, we will prove key lemmata that will be used quite frequently to develop the proof of the main result in Theorem 2. The following lemma is similar to [Chenetal2019, Theorem 8] but also holds when the canonical Markov matrix of $P$ is not reversible.
@@
Let $\pi,{\hat{\pi}},\tilde{\pi}$ be the stationary distributions of the row stochastic matrices $S,{\hat{S}},\tilde{S}$, respectively such that $\pi>0$ entrywise. Then, if $\|\Pi^{\frac{1}{2}}S\Pi^{-\frac{1}{2}}-\sqrt{\pi}\sqrt{\pi}^{\mathrm{T}}\|_{2}+\|S-\hat{S}\|_{\pi^{-1}}<1$, we have
$\|{\hat{\pi}}-\tilde{\pi}\|_{\pi^{-1}}\leq\frac{\|\tilde{\pi}^{\mathrm{T}}(\tilde{S}-{\hat{S}})\|_{\pi^{-1}}}{1-\|\Pi^{\frac{1}{2}}S\Pi^{-\frac{1}{2}}-\sqrt{\pi}\sqrt{\pi}^{\mathrm{T}}\|_{2}-\|S-\hat{S}\|_{\pi^{-1}}}.$
Proof.
By stationarity of $\tilde{S}$ and $^{\hat{S}}$, we have
@@
$\|\tilde{\pi}-{\hat{\pi}}\|_{\pi^{-1}}$ $\leq\|\tilde{\pi}^{\mathrm{T}}(\tilde{S}-{\hat{S}})\|_{\pi^{-1}}+\|\tilde{\pi}-{\hat{\pi}}\|_{\pi^{-1}}\|\hat{S}-S\|_{\pi^{-1}}$
$\ \ \ \ \ +\|\tilde{\pi}-{\hat{\pi}}\|_{\pi^{-1}}\|S-\mathbf{1}_{n}\pi^{\mathrm{T}}\|_{\pi^{-1}}.$ (43)
@@
$\|S-\mathbf{1}_{n}\pi^{\mathrm{T}}\|_{\pi^{-1}}\!=\!\|\Pi^{\frac{1}{2}}S\Pi^{-\frac{1}{2}}-\sqrt{\pi}\sqrt{\pi}^{\mathrm{T}}\|_{2}=\sigma_{2}(\Pi^{\frac{1}{2}}S\Pi^{-\frac{1}{2}}),$
where $\sigma_{2}(M)$ denotes the second-largest singular value of $M$, and the last equality follows since $\sqrt{\pi}$ is both the left and right top singular vector of the DTM $\Pi^{1/2}S\Pi^{-1/2}$ [makur2019, Proposition 2.2]. Thus, rearranging the terms in the above inequality establishes the statement of Lemma 1. ∎
We would like to emphasize the distinction between our proof above and the approach adopted in [Chenetal2019, Theorem 8] for proving a similar result, and elucidate the proof in [Chenetal2019]. In transitioning from Eq. 42 to Eq. 43, the authors in [Chenetal2019] utilize the $\ell^{2}(\pi)$ norm, as opposed to the $\ell^{2}(\pi^{-1})$ norm. Specifically, they bound the final term on the right-hand side as follows
@@
Moreover, since the Markov chain is irreducible by the Perron-Frobenius theorem, we have $\lambda_{2}(S)<1$, and the corresponding upper bounds have been derived in [NegahbanOhShah2012, Chenetal2019]. However, we are interested in the general case where $S$ need not be reversible (as the underlying pairwise comparison matrix may not be BTL). Hence, we bound each of these terms using the following two lemmata.
Lemma 2 (Spectral Norm of Noise).
For $^{\hat{S}}$ constructed as in Eq. 23, we have
@@
The proof of Lemma 2 is quite similar to [NegahbanOhShah2012] but requires slight modifications and is provided in Section V-C3 for completeness.
Next we upper bound the quantity $\|R-\sqrt{\pi}\sqrt{\pi}^{\mathrm{T}}\|_{2}$. Observe that $\|R-\sqrt{\pi}\sqrt{\pi}^{\mathrm{T}}\|_{2}$ is the second largest singular value of the DTM $R$ (as $\sqrt{\pi}$ are both the left and right singular vectors of $R$ corresponding to largest singular value of $1$ [makur2019, Proposition 2.2]). The second largest singular value of DTM $R$ is also equal to the square root of the contraction coefficient for $\chi^{2}$-divergence $\eta_{\chi^{2}}(\pi,{S})$ of the source-channel pair $(\pi,S)$ (see [sarmanov1958, makur2019, MakurZheng2020] for definitions and details). For the case of the complete graph, we can upper bound this quantity by upper bounding the Dobrushin contraction coefficient for TV distance, as demonstrated in the following lemma.
@@
Proof.
We will find an upper bound on $\sqrt{\eta_{\chi^{2}}(\pi,{S})}$ using [makur2019, Proposition 2.5]:
$\eta_{\chi^{2}}(\pi,S)\leq\eta_{\mathsf{TV}}(S)=\max_{i,j\in[n]}\ \left\|S_{i,:}-S_{j,:}\right\|_{\mathsf{TV}},$
@@
$\left\|{S}_{i,:}-{S}_{j,:}\right\|_{\mathsf{TV}}$ $=1-\sum_{k=1}^{n}\min\{S_{ik},S_{jk}\}$
$=1-\left(\sum_{k:k\neq i,j}\frac{\min\{p_{ik},p_{jk}\}}{2n}+\frac{p_{ij}+p_{ji}}{2n}\right)$
@@
Hence, the lemma holds since for $0\leq x\leq 1$, we have $\sqrt{1-x}\leq 1-\frac{x}{2}$. ∎
In the more general case of an arbitrary graph (consistent with our assumptions), we upper-bound the second largest singular value of $R$ by leveraging edge expansion properties and the Cheeger inequality for non-negative matrices, as demonstrated in the following lemma.
@@
Let $\mathcal{G}$} be the induced graph corresponding to the canonical Markov matrix $S$. Consider the DTM $R=\Pi^{1/2}S\Pi^{-1/2}$, with edge expansion of $S$ lower bounded by $\xi$, i.e., $\phi(S)\geq\xi$, then the following bounds hold on the second largest singular value of $R$:
@@
Proof.
Observe that
$\sigma_{2}(R)$ $=\sqrt{\lambda_{2}\left(RR^{\mathrm{T}}\right)}\stackrel{{\scriptstyle\zeta_{1}}}{{\leq}}\sqrt{\lambda_{2}\left(\frac{R+R^{\mathrm{T}}}{2}\right)}$
$\stackrel{{\scriptstyle\zeta_{2}}}{{\leq}}\sqrt{1-\frac{1}{2}\phi^{2}\left(\frac{R+R^{\mathrm{T}}}{2}\right)}$